Stan does 171 J of work lifting himself
0.4 m.
What is Stan’s mass? The acceleration of
gravity is 9.8 m/s2 .
Answer in units of kg
1
Expert's answer
2012-12-24T11:03:33-0500
The applied force is
F = A/S = 171/0.4 = 427.5 N
The mass is
m = F/g = 427.5/9.8 = 43.6 kg